OpenStudy (aaronq):

what does tan2x equal when, x=-60 degrees?

OpenStudy (anonymous):

2*-60 is 120 so cos(-120) = \[\frac{ 1 }{ 2 }\] and sin(-120) = \[\frac{ -\sqrt{3} }{ 2 }\] tan is opp/adj which will then give you an exact answer of \[\sqrt{3}\]

OpenStudy (anonymous):

tan = sin/cos

OpenStudy (aaronq):

thanks so much, i wasn't sure if i could just multiply 2 by 60, or if i had to use the identity 2tanx/1-tan^2x

OpenStudy (anonymous):

as long as the 2 and x are together you are fine tan(2x)

OpenStudy (aaronq):

what if it's tan2x? it didn't necessarily specify tan(2x)

OpenStudy (anonymous):

That is fine also if the 2 was outside for example 2tan(x) then it would be different

OpenStudy (aaronq):

ok sweet, thanks (Y)